OpenStudy (anonymous):

The mean of a set of 5 different positive integers is 21. How many values are possible for the median of this set of positive integers? Please help:)

OpenStudy (anonymous):

This is a hard one. I don't know if this is the right solution, but here's my approach: The 5 integers must be different. So to find the possible values (range) of the middle integer (the median) we set up the other integers like this, starting from 1: 1+2+c+d+e = 105 Then use algebra to reduce: c+d+e = 102 Then set up again using DIFFERENT integers than 1 and 2: 3 + 4 + 95 = 102 So the minimum value for the median number is 3. The equation would look like this: (1+2+3+4+95=105)/5=21 But how to find the max value for the middle integer? I don't know if there's a mathematical way to approach this. I just assumed that since the numbers must be in order of 0<a<b<c<d<e, then we have to set up MAX so that the three numbers equal 102, but they are different than each other and each one is greater than the previous: c<d<e, and c+d+e=102 Doing some guesswork, I know that 33 x 3 = 99, so try this: 33+34+35=102 THat works. So the min value is 3, and the max is 33. I'd be curious if there's a faster way to do this.

OpenStudy (anonymous):

Cool xD

OpenStudy (anonymous):

Endo, why did you set up the integers to equal 105, though?